OpenStudy (anonymous):

evaluate the exponential expression -2^4

OpenStudy (anonymous):

\(-2^4=-2\times 2\times 2\times 2\)

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

yeah, that too!